OpenStudy (anonymous):

I really need help. I am really confused on how to fid the "zeros" of an equation?

OpenStudy (bibby):

Finding the zeroes of an equation usually means setting it (the equation) equal to zero

OpenStudy (bibby):

For example find the zeros of \[x^2+5x+6\] You'd first factor the expression into: \[(x+3)(x+2)\] and then set that equal to zero. You then set each part equal to zero because in a product x*y The resulting product is only 0 if x =0 or y = 0

OpenStudy (anonymous):

x3 – 5x2 + 6x = 0 can you talk me through this equation because that didn't really help me

OpenStudy (bibby):

factor out the x\[x^3 – 5x^2 + 6x = x(x^2 – 5x + 6)\] does that help

OpenStudy (anonymous):

okay yea

OpenStudy (bibby):

thx 4 the medal bby